# Semiparametric Skew-Elliptical Distributions For High-Dimensional Graphical Models

### v1

@@ Abstract
We propose a semiparametric approach called elliptical skew-(S)KEPTIC for efficiently and robustly estimating non-Gaussian graphical models. Relaxing the assumption of semiparametric elliptical distributions to the family of meta skew-elliptical that accommodates a skewness component, we derive a new estimator which is an extension of the SKEPTIC estimator in Liu et al., 2012a , based on semiparametric Gaussian copula graphical models, to the case of skew-elliptical copula graphical models. Theoretically, we demonstrate that the elliptical skew-(S)KEPTIC estimator achieves robust parametric convergence rates in both graph recovery and parameters estimation. We conduct numerical simulations to prove the reliable graph recovery performance of the elliptical skew-(S)KEPTIC estimator. Finally, the new method is applied to the daily log-returns of the stocks of the S&P500 index and shows better interpretability compared to the Gaussian copula graphical models.
Keywords—High-dimensional statistics, undirected graphical models, elliptical distributions, skewness, robust statistics.
@@ 1 Introduction
We consider the problem of estimating high-dimensional undirected graphical models. Given a $p$-dimensional random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$, we aim to estimate the undirected graph $G:=(V,E)$, where the vertices set $V:=\{1,\ldots,p\}$ contains the number of nodes corresponding to the $p$ variables, and the edge set $E$ indicates the conditional independence relationships between $X_{1},\ldots,X_{p}$. Let $X_{i}$ and $X_{j}$ be random variables associated with the multivariate distribution $\boldsymbol{X}$ and define $X_{\setminus{\{i,j}\}}$ the set of variables excluding $X_{i}$ and $X_{j}$. If $X_{i}$ and $X_{j}$ are conditionally independent given $X_{\setminus{\{i,j}\}}$, then there exists no edge between the corresponding nodes $i$ and $j$ in the graph G, denoted as $(i,j)\notin E$.
@@
One reliable approach to estimate the undirected graphical models is the semiparametric nonparanormal (NPN) method proposed in Liu et al., (2009), which relaxes the assumption of Gaussian data, considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that: $f(\boldsymbol{X})=(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim N(0,\Sigma^{0})$, where $\Omega=(\Sigma^{0})^{-1}$. When $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, no edge connects $X_{i}$ and $X_{j}$ if $\Omega_{ij}=0$. Uncovering sparsity is subject to the estimation of the unknown correlation matrix. Liu et al., 2012a propose the SKEPTIC (Spearman/Kendall estimates preempt transformations to infer correlation) estimator for the nonparanormal distribution, a regularized rank-based correlation coefficient estimator based on Kendall’s tau and Spearman’s rho statistics. The estimated correlation matrix is then puggled into shrinkage procedures that reveal the sparsity of the graph.
@@
The nonparanormal distribution exploits a Gaussian-based kernel and belongs to the broad class of meta-elliptical distributions described in Han and Liu, (2014), as well as Liu et al., 2012b , where a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\mathcal{E}C(0,\Sigma^{0},\xi)$ is considered. In particular, Liu et al., 2012b show that the graph $G$ is encoded in the latent generalized partial correlation matrix $\Xi:=-\left[\text{diag}(\Omega)\right]^{-1/2}\Omega\left[\text{diag}(\Omega)\right]^{-1/2}$, thus no edge connects $X_{i}$ and $X_{j}$ if $\Xi_{ij}=0$, which implies $\Omega_{ij}=0$. The main drawback of the meta-elliptical class of distributions is that, even though it embeds skewed distributions, the skewness term does not contribute to reveal the conditional independence structure. For this reason, we introduce the meta skew-elliptical family of distributions defined as $\boldsymbol{X}\sim MS\mathcal{E}(0,\Sigma^{0},\Lambda;f)$ considering a set of functions $f$ such that $f(\boldsymbol{X})\sim SU\mathcal{E}(0,\Sigma^{0},\Lambda,h^{(p+q)},\tau,\Gamma)$ described in Arellano-Valle and Genton, 2010b .
@@
In this paper, we propose a rank-based correlation matrix estimator, which we refer to as elliptical skew-(S)KEPTIC estimator, extending the SKEPTIC estimator to the $MS\mathcal{E}$ family of distributions, adding respect to nonparanormal models the possibility to rely on graphical models that take into account other distributional features such as directional skewness and heavy tails. The main idea is to consider the $\operatorname{SU\mathcal{E}}$ as a combination of the closed skew-Normal distribution introduced in Domìnguez-Molina et al., (2003), which belongs to the class of $\operatorname{SU\mathcal{E}}$ family of distributions, and possesses the stochastic representation defined through Normal distribution and a half-Normal distribution, for which the SKEPTIC estimator of the correlation matrix is defined. The estimated correlation matrix is plugged into existing shrinkage procedures. For instance, Meinshausen and Bühlmann, (2006) study a neighborhood selection approach with the Lasso operator, Banerjee et al., (2008) and Friedman et al., (2008) propose an approach based on penalized likelihood, and Yuan, (2010) and Cai et al., (2011) introduce, respectively, Graphical Dantzig selector and CLIME which are two methods based on linear programming that provide an improvement in terms of theoretical properties respect to the other shrinkage procedures.
@@
We generalize the results in Liu et al., 2012a extending the SKEPTIC estimator to the elliptical skew-(S)KEPTIC estimator by introducing a directional skewness term, encoded in the matrix $\Lambda$, which assigns greater importance to the asymmetry present in data distributions. The proposed estimator is a version of the SKEPTIC estimator conforming to the ${MS}\mathcal{E}$ distribution and embeds two new subclasses of estimators based on Kendall’s tau and Sparman’s rho statistics. While the nonparanormal distribution is a flexible approach to model non-Normal data, the meta skew-elliptical distribution allows us to quantify the asymmetry (and the kurtosis) presented in the data. When the $\operatorname{SU\mathcal{E}}$ distribution reduces to the $\operatorname{CSN}$ distribution, we consider both the statistics for estimating the unknown correlation matrix and we define the elliptical skew-SKEPTIC estimator. Otherwise, we only consider Kendall’s tau statistic, and we define elliptical skew-KEPTIC estimator since Spearman’s rho statistic is not invariant for the class of elliptical distributions (Hult and Lindskog, , 2002).
The proposed framework extends to a greater variety of semiparametric graphical models, introducing the possibility of modeling the combination of directional skewness and heavy tails data and adding flexibility to the nonparanormal-type of data transformations.
We investigate the theoretical properties of the elliptical skew-(S)KEPTIC estimator, and find bounds equal to twice the bounds of the SKEPTIC estimator. The increase in model flexibility justifies the slight reduction in the robustness of the proposed estimators.
@@
We provide numerical studies to support our theory. Through a backward approach, we define the true sparsity pattern given by $\Omega$ and provide the graph recovery using the elliptical skew-(S)KEPTIC estimator which is compared to the SKEPTIC estimator. The elliptical skew-(S)KEPTIC estimator proves an optimal graph recovery, giving statistical efficiency to the proposed method.
@@
Prior studies by Zareifard et al., (2016), Nghiem et al., (2022), and Sheng et al., (2023) have explored graphical models while incorporating skewness into the sparsity pattern estimation. Zareifard et al., (2016) exploit a multivariate $\operatorname{CSN}$ distribution to define the skew Gaussian graphical models which are estimated through a Bayesian approach. Nghiem et al., (2022) propose a novel nodewise regression approach to estimate the graph on data generated from a generalized multivariate skew-Normal distribution. Instead, Sheng et al., (2023) include the shape parameter in the determination of the precision matrix using the skew-Normal distribution of Azzalini, (1985). They build an algorithm that penalizes the likelihood of this distribution in order to estimate $\Omega$ and thus the graph. However, these studies do not emphasize the direct estimation of the correlation matrix while accounting for skewness.
The paper is organized as follows: in Section 2 we recall the background, in Section 3 we define the ${MS}\mathcal{E}$ distribution, which we use to derive the elliptical skew-(S)KEPTIC estimator and in Section 4 we study the theoretical properties of the proposed estimator. In Section 5, we conduct experiments comparing elliptical skew-(S)KEPTIC estimator with the SKEPTIC estimator through numerical comparisons on synthetic data and with an empirical analysis on the S&P 500 stocks data. Section 6 concludes.
@@ 2 Background
In this section, we briefly recall the literature preceding our article: the SKEPTIC estimator of Liu et al., 2012a , the semiparametric ellitpical distributions in Liu et al., 2012b and Han and Liu, (2014), in this work, we rely of the meta-elliptical distribution.. The section concludes with the introduction of $\operatorname{SU\mathcal{E}}$ family of distributions in Arellano-Valle and Genton, 2010b , and its link with the undirected graphical models.
@@ 2.1 Notation
Here is the notation used throughout the paper. Let $A=\left[A_{ij}\right]\in\mathbb{R}^{p\times p}$ and $\boldsymbol{a}=(a_{1},\ldots,a_{p})\in\mathbb{R}^{p}$. For $1\leq q<\infty$, we define $\|\boldsymbol{a}\|_{q}=\left(\sum_{i=1}^{p}|\boldsymbol{a}|^{q}\right)^{1/q}$, when $q=\infty$: $\|\boldsymbol{a}\|_{\infty}=\max_{1\leq i\leq p}|\boldsymbol{a}_{i}|$. Regarding the $A$ matrix, when $q=1$: $\|A\|_{1}=\max_{1\leq j\leq q}\sum_{i=1}^{p}|A_{ij}|$, and when $q=\infty$, $|A|_{\infty}=\max_{1\leq i\leq q}\sum_{j=1}^{p}|A_{ij}|$. We denote $\|A\|_{\max}=\max_{i,j}|A_{ij}|$. A vector without its $j$-th element is defined: $\boldsymbol{a}_{\backslash j}=(\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{j-1},\boldsymbol{a}_{j+1},\ldots,\boldsymbol{a}_{p})^{\top}\in\mathbb{R}^{p-1}$, and we denote $A_{\backslash i,\backslash j}\in\mathbb{R}^{p-1\times p-1}$ the matrix without the $i$-$th$ row and the $j$-th column. We define: $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ respectively the minimum and the maximum eigenvalue of $A$. $\Omega$ denotes a sparse matrix, and we define the degree of $\Omega$: $\operatorname{deg}(\Omega)=\max_{1\leq i\leq p}\sum_{j=1}^{p}\operatorname{I}(|\Omega_{ij}|\neq 0)$.
@@ 2.2 The Nonparanormal SKEPTIC
A random vector $\boldsymbol{X}=(X_{1},...,X_{p})^{\top}$ has a nonparanormal distribution if there exists functions $\{f_{j}\}_{j=1}^{p}$ such that $\boldsymbol{Z}\equiv f(\boldsymbol{X})\sim\operatorname{N}(0,\Sigma^{0})$, where $f(\boldsymbol{X})=(f_{1}(X_{1}),...,f_{p}(X_{p}))^{\top}$, we then write $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$. Further, if the functions $\{f_{j}\}_{j=1}^{p}$ are monotone and differentiable, the joint distribution of $\boldsymbol{X}$ is a Gaussian copula. The focal point is to find a function, $f(\boldsymbol{X})$, that transforms our data into normally distributed data. Suppose we have $n$ observations for the $j$-th variable, $x_{1j},...,x_{nj}$, we compute the rank of each observation, and indicate the rank with $r_{1j},...,r_{nj}$. Let $\Phi(\cdot)$ denote the Gaussian cumulative distribution function, the $f(\boldsymbol{X})$ that transforms the data is:
$\hat{f}_{j}(x_{ij})=\Phi^{-1}(\hat{r}_{ij}),\quad\text{where:}\quad\hat{r}_{ij}=\frac{r_{ij}}{n+1}.$ (1)
We consider the following statistics:

### v2

# Robust Semiparametric Graphical Models with Skew-Elliptical Distributions

@@ Abstract
We propose semiparametric estimators called elliptical skew-(S)KEPTIC for efficiently and robustly estimating non-Gaussian graphical models, relaxing the assumption of semiparametric elliptical distributions to the family of meta skew-elliptical, which accommodates a skewness component. Theoretically, we demonstrate that the elliptical skew-(S)KEPTIC estimators achieve robust convergence rates in both graph recovery and parameters estimation. We conduct numerical simulations that prove the reliable graph recovery performance of the elliptical skew-(S)KEPTIC estimators. Finally, the new method is applied to the daily log-returns of the stocks of the S&P500 index and shows higher sparsity compared to the Gaussian copula graphical models.
Keywords: Elliptical distributions, robust statistics, skewness, and undirected graphical models.
@@ 1 Introduction
We consider the problem of estimating undirected graphical models. Given a $p$-dimensional random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$, we aim to estimate the undirected graph $G:=(V,E)$, where the vertices set $V:=\{1,\ldots,p\}$ contains the number of nodes corresponding to the $p$ variables, and the edge set $E$ indicates the conditional independence relationships between $X_{1},\ldots,X_{p}$. Let $X_{i}$ and $X_{j}$ be random variables associated with the multivariate distribution $\boldsymbol{X}$ and define $X_{\setminus{\{i,j}\}}$ the set of variables excluding $X_{i}$ and $X_{j}$. If $X_{i}$ and $X_{j}$ are conditionally independent given $X_{\setminus{\{i,j}\}}$, then there exists no edge between the corresponding nodes $i$ and $j$ in the graph G, denoted as $(i,j)\notin E$.
@@
One reliable approach to estimate the undirected graphical models is the semiparametric nonparanormal (NPN) method proposed in Liu et al., (2009), which relaxes the assumption of Gaussian data, considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that: $f(\boldsymbol{X})=(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{N}(0,\Sigma^{0})$, where $\Omega=(\Sigma^{0})^{-1}$. When $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, no edge connects $X_{i}$ and $X_{j}$ if $\Omega_{ij}=0$. Uncovering sparsity is subject to the estimation of the unknown correlation matrix. Liu et al., 2012a propose the SKEPTIC (Spearman/Kendall estimates preempt transformations to infer correlation) estimator for the nonparanormal distribution, a regularized rank-based correlation coefficient estimator based on Kendall’s tau and Spearman’s rho statistics. The estimated correlation matrix is then plugged into shrinkage procedures that reveal the sparsity of the graph.
@@
The nonparanormal distribution exploits a Gaussian-based kernel and belongs to the broad class of meta-elliptical distributions described in Han and Liu, (2014) and in Liu et al., 2012b , where a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\mathcal{E}C(0,\Sigma^{0},\xi)$ is considered. In particular, Liu et al., 2012b show that the graph $G$ is encoded in the latent partial correlation matrix $\Xi:=-\left[\text{diag}(\Omega)\right]^{-1/2}\Omega\left[\text{diag}(\Omega)\right]^{-1/2}$, thus no edge connects $X_{i}$ and $X_{j}$ if $\Xi_{ij}=0$, which implies $\Omega_{ij}=0$. However, as shown by Rossell and Zwiernik, (2021), the dependence structure of the elliptical distribution differs from that of the Gaussian case. In addition, a major pitfall of the meta-elliptical class of distributions is that, even though it embeds skewed distributions, the skewness term does not contribute to reveal the conditional independence structure.
We propose rank-based correlation matrix estimators, which we refer to as elliptical skew-(S)KEPTIC estimators. These estimators refer to the second moment of the meta skew-elliptical family of distributions, that we introduce in this paper considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\operatorname{SU}\mathcal{E}(0,\Sigma^{0},\Lambda,h^{(p+q)},\tau,\Gamma)$, where the unified skew-elliptical ($\operatorname{SU}\mathcal{E}$) distribution is described in Arellano-Valle and Genton, 2010b . Extending the SKEPTIC estimator to the meta skew-elliptical family of distributions, our methodology allows for a semiparametric estimation of graphical models that take into account distributional features such as directional skewness and heavy tails. The main idea is to consider the unified skew-elliptical distribution as a combination of the closed skew-normal distribution introduced in Domìnguez-Molina et al., (2003), which belongs to the class of $\operatorname{SU\mathcal{E}}$ distributions. The class of unified skew-elliptical distributions, which generalizes the closed skew-normal, enjoys key closure properties under marginalization, conditioning, and linear transformations. These properties enable e.g., to define conjugate priors in probit, tobit, and multinomial probit regression models based on $\operatorname{SU\mathcal{E}}$ distributions, as both the prior and posterior remain within the same family (Durante,, 2019, Karling et al.,, 2024).
@@
The unified skew-elliptical distribution can be written in terms of the closed-skew normal, that possesses a stochastic representation defined through a normal distribution and a half-normal distribution, for which the SKEPTIC estimator of the correlation matrix is defined. The estimated correlation matrix is then plugged into existing shrinkage procedures. For instance, Meinshausen and Bühlmann, (2006) study a neighborhood selection approach with the lasso operator, Banerjee et al., (2008) and Friedman et al., (2008) propose an approach based on penalized likelihood, and Yuan, (2010) and Cai et al., (2011) introduce, respectively, graphical Dantzig selector and CLIME which are two methods based on linear programming that provide an improvement in terms of theoretical properties respect to the other shrinkage procedures.
@@
The proposed estimators generalize the SKEPTIC estimator conforming to the meta skew-elliptical distribution and embed two new subclasses of estimators based on Kendall’s tau and Spearman’s rho statistics. While the nonparanormal distribution is a flexible approach to model non-Gaussian data, the approach based on the meta skew-elliptical distribution allows us to quantify the asymmetry and the kurtosis presented in the data. The elliptical skew-(S)KEPTIC estimators possess skewness terms, encoded in the matrix of skewness parameters $\Lambda$, which assigns greater importance to the asymmetry present in data distributions. When the unified skew-elliptical distribution reduces to the closed skew-normal distribution, we consider both the statistics for estimating the unknown correlation matrix and we define the elliptical skew-SKEPTIC estimator. Otherwise, we only consider Kendall’s tau statistic and we define the elliptical skew-KEPTIC estimator, since Spearman’s rho statistic is not invariant for the class of elliptical distributions (Hult and Lindskog,, 2002). The proposed framework extends to graphical models that incorporate both skewness and heavy tails, adding flexibility to nonparanormal-type data transformations.
@@
We prove that the elliptical skew-(S)KEPTIC estimators have deviation from the true correlation matrix $\Sigma^{0}$ that decreases as the sample size grows, showing a fast exponential rate of convergence in $\|\cdot\|_{\max}$ norm. The numerical analysis supports the theory. Through a backward approach, we define the true sparsity pattern given by $\Omega$ and show the graph recovery using the elliptical skew-(S)KEPTIC estimators, which are compared to the SKEPTIC estimator. The elliptical skew-(S)KEPTIC estimators prove an optimal graph recovery, showing statistical efficiency of the proposed method.
@@
Prior studies by Zareifard et al., (2016), Nghiem et al., (2022), and Sheng et al., (2023) have explored graphical models while incorporating skewness into the sparsity pattern estimation. Zareifard et al., (2016) exploit a multivariate closed skew-normal distribution to define the skew Gaussian graphical models, which are estimated through a Bayesian approach. Nghiem et al., (2022) propose a novel nodewise regression approach to estimate the graph on data generated from a generalized multivariate skew-normal distribution. Instead, Sheng et al., (2023) include the shape parameter in the determination of the precision matrix using the skew-normal distribution of Azzalini, (1985). They develop an algorithm that penalizes the likelihood of this distribution to estimate $\Omega$ and, consequently, the graph. However, these studies do not focus on the direct estimation of the correlation matrix while incorporating skewness.
The paper is organized as follows. In Section 2 we describe the background, including a review of the SKEPTIC estimator and the semiparametric elliptical distribution. In Section 3 we introduce the meta skew-elliptical distribution, which forms the basis to derive the elliptical skew-(S)KEPTIC estimators, whereas in Section 4 we provide some shrinkage procedures to estimate the precision matrix $\Omega$. Theoretical properties of the proposed estimator are discussed in Section 5. Section 6 presents simulation analysis comparing the elliptical skew-(S)KEPTIC estimators with the SKEPTIC estimator on synthetic data, and reports an empirical application to S&P 500 stock returns. Section 7 concludes the paper.
@@ 2.1 Notation
Here is the notation used throughout the paper. Let $A=\left[A_{ij}\right]\in\mathbb{R}^{p\times p}$ and $\boldsymbol{a}=(a_{1},\ldots,a_{p})\in\mathbb{R}^{p}$. For $1\leq q<\infty$, we define $\|\boldsymbol{a}\|_{q}=\left(\sum_{i=1}^{p}|\boldsymbol{a}|^{q}\right)^{1/q}$, when $q=\infty$: $\|\boldsymbol{a}\|_{\infty}=\max_{1\leq i\leq p}|\boldsymbol{a}_{i}|$. Regarding the $A$ matrix, when $q=1$: $\|A\|_{1}=\max_{1\leq j\leq q}\sum_{i=1}^{p}|A_{ij}|$, and when $q=\infty$, $\|A\|_{\infty}=\max_{1\leq i\leq q}\sum_{j=1}^{p}|A_{ij}|$. We denote $\|A\|_{\max}=\max_{i,j}|A_{ij}|$. A vector without its $j$-th element is defined: $\boldsymbol{a}_{\backslash j}=(\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{j-1},\boldsymbol{a}_{j+1},\ldots,\boldsymbol{a}_{p})^{\top}\in\mathbb{R}^{p-1}$, and we denote $A_{\backslash i,\backslash j}\in\mathbb{R}^{p-1\times p-1}$ the matrix without the $i$-thth row and the $j$-th column. We define: $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ respectively the minimum and the maximum eigenvalue of $A$. $\Omega$ denotes a sparse matrix, and we define the degree of $\Omega$: $\operatorname{deg}(\Omega)=\max_{1\leq i\leq p}\sum_{j=1}^{p}\operatorname{I}(|\Omega_{ij}|\neq 0)$.
@@ 2.2 The Nonparanormal SKEPTIC
A random vector $\boldsymbol{X}=(X_{1},...,X_{p})^{\top}$ has a nonparanormal distribution if there exists functions $\{f_{j}\}_{j=1}^{p}$ such that $\boldsymbol{Z}\equiv f(\boldsymbol{X})\sim\operatorname{N}(0,\Sigma^{0})$, where $f(\boldsymbol{X})=(f_{1}(X_{1}),...,f_{p}(X_{p}))^{\top}$, we then write $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$. Further, if the functions $\{f_{j}\}_{j=1}^{p}$ are monotone and differentiable, the joint distribution of $\boldsymbol{X}$ is a Gaussian copula. The focal point is to find a function, $f(\boldsymbol{X})$, that transforms our data into normally distributed data.
Given $n$ observations of the $j$-th variable, $x_{1j},\dots,x_{nj}$, we define the empirical cumulative distribution function (ECDF) $\hat{F}_{j}(w)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(x_{ij}\leq w),$ where $\mathbb{I}(\cdot)$ is the indicator function. The transformation function $\hat{f}_{j}(x_{ij})$ is then defined as $\hat{f}_{j}(x_{ij})=\Phi^{-1}(\hat{F}_{j}(x_{ij}))$, where $\Phi^{-1}(\cdot)$ denotes the quantile normal distribution function. In practice, the ECDF at each observation $x_{ij}$ can be approximated using its normalize rank: $\hat{F}_{j}(x_{ij})\approx\hat{r}_{ij}=\frac{r_{ij}}{n+1}$, where $r_{ij}$ is the rank of $x_{ij}$ among $x_{1j},\dots,x_{nj}$. Thus, the transformation results in
$\hat{f}_{j}(x_{ij})=\Phi^{-1}\left(\hat{r}_{ij}\right).$ (1)
This approach ensures that the transformed data approximate a standard normal distribution while preserving the rank structure of the original data. We consider the following statistics:
$\text{(Spearman's rho)}\ \hat{\rho}_{ij}=\frac{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})(r^{k}_{i}-\bar{r}_{i})}{\sqrt{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})^{2}\sum_{k=1}^{n}(r^{k}_{i}-\bar{r}_{i})^{2}}},$ (2)
$\text{(Kendall's tau)}\ \hat{\tau}_{ij}=\frac{2}{n(n-1)}\sum_{1\leq k<k^{\prime}\leq n}\text{sign}((x^{k}_{j}-x^{k^{\prime}}_{j})(x^{k}_{i}-x^{k^{\prime}}_{i})).$ (3)
Both $\hat{\rho}_{ij}$ and $\hat{\tau}_{ij}$ are non-parametric correlations between the empirical realizations of random variables $X_{j}$ and $X_{i}$. Let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, $F_{i}$ and $F_{j}$ denote the cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$, the population versions of Spearman’s rho and Kendall’s tau are defined as: $\rho_{ij}:=\operatorname{Corr}(F_{i}(X_{i}),F_{j}(X_{i}))$ and $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$, respectively. For nonparanormal distributions, there is a connection between Spearman’s rho and Kendall’s tau to the underlying Pearson’s correlation coefficient $\Sigma^{0}_{ij}$. We recall the SKEPTIC estimator.
Lemma 2.1.
(Kruskal, (1958), Liu et al., 2012a ): Assuming $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, we have $\Sigma_{ij}^{0}=2\operatorname{sin}\left(\frac{\pi}{6}\rho_{ij}\right)=\operatorname{sin}\left(\frac{\pi}{2}\tau_{ij}\right)$.
From these results two estimators of the unknown correlation matrix $(\Sigma^{0}_{ij})$ are $\hat{S}^{\rho}=\left[\hat{S}^{\rho}_{ij}\right]$ and $\hat{S}^{\tau}=\left[\hat{S}^{\tau}_{ij}\right]$, where: $\hat{S}^{\rho}_{ij}=2\operatorname{sin}\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ and $\hat{S}^{\tau}_{ij}=\operatorname{sin}\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$. Spearman’s rho and Kendall’s tau statistics are invariant under strictly increasing transformations of the underlying random variable.
2.3 The meta-elliptical distribution
The meta-elliptical distribution introduced in Han and Liu, (2014) relaxes the assumption of elliptically distributed data, in a similar way as the nonparanormal of Liu et al., (2009) does for Gaussian distributions. The meta-elliptical family is a strict extension of the nonparanormal distribution.
Definition 2.2.
(Han and Liu, (2014), Fang, (2018)) A random vector $\boldsymbol{X}$ has an elliptical distribution, $\boldsymbol{X}\sim\mathcal{E}C_{p}(\mu,\Sigma,\xi)$, if $\boldsymbol{X}$ has the following stochastic representation: $\boldsymbol{X}\overset{d}{=}\mu+\xi AU$. Where: $\mu\in\mathbb{R}^{p}$, $q:=rank(A)$, AA is a $p\times q$ matrix such that $\Sigma=AA^{\top}$, $\xi\geq 0$ is a random variable independent of $U$, $U\in\mathbb{S}^{q-1}$, which is uniformly distributed of the unit sphere in $\mathbb{R}^{q}$.
Definition 2.3.
(Han and Liu, (2014)) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has a meta-elliptical distribution, denoted by $\boldsymbol{X}\sim M\mathcal{E}_{p}(\Sigma^{0},\xi;f)$, if there exist a set of univariate monotone functions $f_{1},\ldots,f_{p}$ such that $(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\mathcal{EC}_{p}(0,\Sigma^{0},\xi)$, where $\Sigma^{0}\in\mathcal{R}_{p}$ is the latent correlation matrix.
For the distributions that belong to the meta-elliptical family, Kendall’s tau statistic in Lemma 2.1 can be exploited to estimate the unknown correlation matrix. Nevertheless, the meaning of a missing edge is different for the nonparanormal case. Indeed, when $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$ a missing edge implies conditional independence. On the other hand, when $\boldsymbol{X}\sim M\mathcal{E}(\Sigma^{0},\xi,f)$ a missing edge means uncorrelation.